# Ada-VE: Training-Free Consistent Video Editing Using Adaptive Motion Prior

### v1

@@ 4.1.1 Successive Motion Mask Extraction
Following prior work [42, 11, 44, 30], we initially sample a total of $Z$ reference frames $\mathbf{R}=\{R_{1},R_{2},\dots,R_{Z}\}$ from a total of $N$ frames for joint-editing. The sampling interval $s$ can be adjusted based on the video length and computational overhead. Instead of naively extending cross-frame attention across $Z$ reference frames with inflated T2I diffusion models [44, 11], we propose adaptive sparse extension of KV features to reduce redundant computations. We extract successive motion masks $M_{i}$ from these $Z$ frames to separate moving regions from stationary regions, which are later used for sparse extended self-attention operations (see Fig. 3).
@@
${F}_{i}=\mathtt{FlowEst}(R_{i-1}\rightarrow R_{i}),\ G_{i}=\mathtt{RGB2Gray}\left(\mathtt{Flow2RGB}(F_{i})\right),\ \forall i=\{2,3,\dots,Z\}$ (3)
Afterwards, stored gray-scale optical flow visualizations $G_{i}$ of each reference frame are used in every self-attention block for sparse sampling of KV features. Each self-attention block operates at a smaller spatial resolution than $(h\times w)$. However, LDM preserves location features throughout the network, facilitating the reuse of spatial motion priors encoded in $G_{i}$. Hence, we perform downsampling on $G_{i}$ to match the spatial resolution $(d^{j}\times d^{j})$ of the feature map $X^{j}$. A simple adaptive thresholding operation (e.g., Otsu’s thresholding [29]) extracts spatial moving region masks $M^{j}$, given by:
$M_{i}^{j}=\mathtt{ApplyThreshold}(\mathtt{DownSample}(G_{i},\tau)),\ \forall i=\{2,\dots,Z\}$ (4)
where $\tau=d^{j}/h$ represents the downsampling ratio at the $j^{th}$ self-attention block. These motion masks are used in the sparse extension of joint self-attention of reference frames.
@@ 4.1.2 Adaptive Extension of Self-Attention
The key idea of adaptive extension of self-attention is to integrate more features of the moving regions while eliminating redundant features from the stationary ones. We use estimated motion masks $M_{i}^{j}$ for aggregating moving region features at every $j^{th}$ self-attention block. However, successive motion masks may lose some key features, especially at the surroundings due to occlusions.
Hence, we use the full-frame KV features at a sparser interval $r$ ($r>s$), and only use the moving region features from other reference frames based on motion mask $M$ (see Fig. 4). The proposed extended sparse KV can be represented as:
$\mathbf{K}_{\mathtt{sparse}}^{j}=\{K_{1}^{j},\dots,(M_{r-1}^{j}\cdot K_{r-1}^{j}),K_{r}^{j},(M_{r+1}^{j}\cdot K_{r+1}^{j}),\dots,K_{2r}^{j},\dots,K_{Z}^{j}\}$ (5)
$\mathbf{V}_{\mathtt{sparse}}^{j}=\{V_{1}^{j},\dots,(M_{r-1}^{j}\cdot V_{r-1}^{j}),V_{r}^{j},(M_{r+1}^{j}\cdot V_{r+1}^{j}),\dots,V_{2r}^{j},\dots,V_{Z}^{j}\}$ (6)
@@
$\mathtt{SparseExtendedSelfAttention}(X_{i})=\mathtt{Softmax}\left(\frac{Q_{i}K_{\mathtt{sparse}}^{T}}{\sqrt{d}}\right)V_{\mathtt{sparse}},\ \text{where}\ I_{i}\in\mathbf{R}$ (7)
The proposed KV-extension in $\mathtt{SESA}(\cdot)$ selectively retains moving region tokens from all frames. Hence, slower motion regions will adaptively incorporate sparser sampling of moving regions, while faster motion regions will leverage denser sampling of KVs.
@@ 4.2 KV-Caching and Inter-Frame Editing
Generating longer-duration videos by considering all frames in joint attention is computationally intensive. To address this, we employ sparser sampling rates of reference frames (higher $s$, $r$) in the initial phase for longer videos. Subsequently, we generate intermediate frames in the second phase. Previous approaches often rely on frame interpolation techniques for intermediate frames [42, 21], but these methods can result in blurring effects or inconsistent character generation. Instead of interpolation, Ada-VE focuses on directly generating intermediate frames with consistent characters to better preserve visual quality.
Our empirical observations indicate that maintaining a consistent KV feature space is crucial for generating consistent characters across intermediate frames. Rather than using only nearby KVs for intermediate frame generation, we propose caching the sparsely sampled KVs from all reference frames (i.e., $\mathbf{K}_{\mathtt{sparse}}$, $\mathbf{V}_{\mathtt{sparse}}$) during the first pass. These cached KVs are then reused with the queries $Q_{m}$ from all intermediate frames $I_{m}$ in self-attention ($\mathtt{IFSA}(\cdot)$) during the second pass:
$\mathtt{InterFrameSelfAttention}(X_{m})=\mathtt{Softmax}\left(\frac{Q_{m}K_{\mathtt{sparse}}^{T}}{\sqrt{d}}\right)V_{\mathtt{sparse}},\ \text{where}\ I_{m}\notin\mathbf{R}$ (8)
@@ 5.1 Main results
We used 15 full-length DAVIS [31] videos and 50 manually designed prompts to conduct quantitative evaluations based on user preference. Users were asked to select the best-edited videos in three categories: (1) Visual Quality: This considers the amount of blurriness and overall visual perception of users on edited videos; (2) Temporal Consistency: This represents the temporal motion smoothness of edited videos; and (3) Content Consistency: This denotes the generation of consistent characters across all frames of each video.
Table 1: User preference study is conducted on visual quality, temporal conistency and content consistency with the same prompts. * denotes shorter video length for memory requirements.
@@
T2V-zero* [18] 1.7 0.33 0.04
SDEdit [25] 3.4 15.81 6.52
ControlVideo [44] 12.2 5.16 1.28
TokenFlow [11] 22.3 40.2 40.74
Ada-VE(ours) 60.4 38.5 51.42
Four baseline methods were used for comparative analysis: Text2Video-zero [18], SDEdit [25], TokenFlow [11], and ControlVideo [44]. Note that only 16 frames of generated videos were considered for Text2Video-zero due to extensive memory requirements for longer-duration videos. Moreover, we considered TokenFlow integrated SDEdit [25] to make it a stronger video baseline for a fair comparison.
The results are presented in Table 1. Ada-VE achieved the highest user preference in visual quality at 67.4%, enhancing visual representation with reduced blurriness, and the highest content consistency at 50.56%, generating consistent content across frames. In contrast, most other video synthesis methods struggled to generate consistent content and often exhibited temporal blurriness. Moreover, Ada-VE performed on par with the state-of-the-art TokenFlow method in terms of temporal consistency, while significantly outperforming other baseline methods. Hence, Ada-VE significantly improves visual quality and content consistency without sacrificing motion smoothness.
@@
We present a qualitative analysis of the baseline methods in Fig. 5. We exclude Text2Video-zero from this analysis due to its limitations in synthesizing longer videos. In the example of the woman running, SDEdit [25] and TokenFlow [11] often produce blurry fingers and hands, missing important details. Additionally, SDEdit generates inconsistent legs and distorted face structures in the moonwalk videos. TokenFlow exhibits similar face distortions. ControlVideo generates inconsistent hairstyles and hands, with more prominent inter-frame inconsistencies in the moonwalk videos. In contrast, Ada-VE generates consistent fingers in the woman running example with significantly reduced motion blurriness. Moreover, the face distortion in the moonwalk videos is notably improved in Ada-VE. These results demonstrate the superior quality of Ada-VE compared to state-of-the-art methods. Additionally, we provide extensive qualitative videos in the supplementary.
Figure 5: Qualitative comparisons across state-of-the-art video editing methods. SDEdit [25] suffers from motion blurring in longer duration videos and inconsistent character generation. ControlVideo also exhibits issues with inconsistent character generation. TokenFlow shows better temporal consistency but suffers from motion blurring. Ada-VE (ours) achieves the best visual quality while maintaining consistency throughout the videos.
Figure 6: We present ablation study on video editing and style transfer prompts.
@@ 5.2 Ablation Study
We study the effects of various extensions of self-attention performance in joint editing, as shown in Table 2. We selected 40 consecutive frames from 10 DAVIS videos for joint editing based on text prompts. To quantitatively measure temporal consistency and realistic prompt alignment, we used the warping-error and CLIP-score metrics following prior work [11, 44]. Apart from the various self-attention extensions, the same baseline PnP model [38] was used for this study. Our observations based on Tab. 2 indicate that increasing the number of reference frames directly enhances performance, although it significantly increases computational burden and latency. Incorporating our proposed adaptive extension of self-attention can significantly reduce latency without compromising performance. For instance, with full extensions of self-attention, Ada-VE can achieve 3.9$\times$ speed-up on average. We use single A6000 GPUs with 48GB memory for experiments.
@@
6 Broader Impact
Ada-VE has the potential to significantly enhance the quality and consistency of video editing across various fields. In the creative industries, it can empower filmmakers, animators, and artists to produce visually stunning content with greater efficiency, leading to more innovative storytelling and artistic expression. Additionally, Ada-VE’s advancements can improve educational and training resources by providing more realistic and engaging visual content, enhancing learning experiences. By making advanced video editing techniques more accessible, Ada-VE enables independent creators and small studios to compete with larger production houses, democratizing high-quality content creation. However, it is important to consider and mitigate potential misuse in areas such as deepfake creation and unauthorized content modification.
@@ 8 Conclusion
In this paper, we presented Ada-VE, an advanced video-to-video synthesis method designed to address the limitations of existing techniques. By introducing an adaptive motion-based KV selection method and the KV-caching technique, Ada-VE effectively reduces computational burden while enhancing visual quality and content consistency. Our approach builds on the Plug-and-Play diffusion models, incorporating sparse KV feature extraction and caching mechanisms to achieve superior performance. Through extensive qualitative and quantitative evaluations, Ada-VE demonstrated significant improvements over state-of-the-art methods, particularly in maintaining visual clarity, temporal consistency, and content fidelity across frames. Our method’s ability to consistently modify moving subjects while preserving fine details was validated in various challenging style transfer and video editing tasks. Overall, Ada-VE represents a substantial advancement in video synthesis, offering a robust solution for generating high-quality, temporally consistent videos. We believe this work opens new avenues for future research in video editing and synthesis, encouraging further exploration and refinement of the proposed methods.
@@ Acknowledgement
This project is supported in part by the ONR Minerva program, iMAGiNE - the Intelligent Machine Engineering Consortium at UT Austin, and a UT Cockrell School of Engineering Doctoral Fellowship.
Appendix A Appendix / supplemental material
A.1 Limitations: Qualitative visualizations
We note that our method primarily built on top of the PnP [38] image editing framework. PnP leverages structural guidance from the given video through extensive feature injection, which often fails to significantly alter the shape of subjects. We present a sample visualization of such failure cases in Fig. 7. It can be seen that the error mostly occurs when attempting to modify a Jeep car into a Porsche car due to their significant shape differences. These issues are largely inherited from the baseline PnP image editing method. Despite this, Ada-VE maintains consistent character across all frames, demonstrating its robustness in extending image models for video editing applications. Additionally, Ada-VE can be easily adapted to any image editing method due to its simple and general architecture.
Figure 7: Visualizing failure cases: Since our method is built on top of Plug-and-Play [38]
diffusion, it cannot inherently follow the modified prompts to change the structure of the subject. Nevertheless, our key contributions are easy to adapt in most exisiting video editing baselines.
Figure 8: Qualitative visualizations of various self-attention extensions: Increasing the number of frames in the extended self-attention significantly improves performance but comes with an extensive computational burden. Ada-VE significantly reduces the operational latency of full extensions while preserving visual quality by leveraging the motion prior of the guidance video. This allows for a substantial increase in the number of frames in joint editing, enabling efficient operation on longer duration videos. For this study, joint editing of a total of 40 frames was used, and we present a portion of these.
A.2 Various extensions of Self-Attention: Qualitative visualizations
We present sample qualitative visualizations of various self-attention extension mechanisms in Fig. 8. We observe consistent character generation when using KVs from a fixed set of frames across the videos. For instance, even when using only the first frame’s KVs, we observe consistent character generation, but the visual quality deteriorates significantly, resulting in structural deformations. Integrating KVs from the immediate previous frames along with the first frame shows some improvement in structural deformations. However, using different KVs across the video leads to significant flickering and inconsistent character generation. For example, the woman’s face is altered, and the background grass is no longer visible.
Using fully extended self-attention significantly improves visual quality and character consistency but increases latency by approximately 13 times due to the computational burden. We observe repeated features across frames, such as the background scenarios of the running woman, which consume a large portion of redundant computation. By integrating Ada-VE, we leverage the motion prior of the guidance video to drastically reduce computational overhead. Ada-VE achieves around a 4$\times$ speed-up for joint editing across these frames. Therefore, Ada-VE can potentially integrate a significantly larger number of frames in joint editing while using similar computational resources.

### v2

@@ 4.1.1 Successive Motion Mask Extraction
Following prior work [44, 11, 47, 32], we initially sample a total of $Z$ reference frames $\mathbf{R}=\{R_{1},R_{2},\dots,R_{Z}\}$ from a total of $N$ frames for joint-editing. The sampling interval $s$ can be adjusted based on the video length and computational overhead. Instead of naïvely extending cross-frame attention across $Z$ reference frames with inflated T2I diffusion models [47, 11], we propose adaptive sparse extension of KV features to reduce redundant computations. We extract successive motion masks $M_{i}$ from these $Z$ frames to separate moving regions from stationary regions, which are later used for sparse extended self-attention operations (see Fig. 3).
@@
$F_{i}$ $=\mathtt{FlowEst}(R_{i-1}\rightarrow R_{i}),\quad\forall i\in\{2,3,\dots,Z\},$ (2)
$G_{i}$ $=\mathtt{RGB2Gray}\left(\mathtt{Flow2RGB}(F_{i})\right)$
Afterwards, stored gray-scale optical flow visualizations $G_{i}$ of each reference frame are used in every self-attention block for sparse sampling of KV features. Each self-attention block operates at a smaller spatial resolution than $(h\times w)$. However, LDM preserves location features throughout the network, facilitating the reuse of spatial motion priors encoded in $G_{i}$. Hence, we perform downsampling (DS) on $G_{i}$ to match the spatial resolution $(d^{j}\times d^{j})$ of the feature map $X^{j}$. A simple adaptive thresholding operation (e.g., Otsu’s thresholding [31]) extracts spatial moving region masks $M^{j}$, given by:
$M_{i}^{j}=\mathtt{ApplyThr}(\mathtt{DW}(G_{i},\tau)),\quad\forall i=\{2,\dots,Z\}$ (3)
where $\tau=d^{j}/h$ represents the downsampling ratio at the $j^{th}$ self-attention block. These motion masks are used in the sparse extension of joint self-attention of reference frames, in every diffusion sampling iteration.
@@ 4.1.2 Adaptive Extension of Self-Attention
The key idea of adaptive extension of self-attention is to integrate more features of the moving regions while eliminating redundant features from the stationary ones. We use estimated motion masks $M_{i}^{j}$ for aggregating moving region features at every $j^{th}$ self-attention block. However, successive motion masks may lose some key features, especially at the surroundings due to occlusions, motion blur, and background movements. To overcome this, we integrate full-frame KV features at a sparser frame interval $r$ ($r>>s$) that overcomes the feature loss. Moreover, we only use the moving region features from most other reference frames based on motion mask $M$ eliminating redundant computations (see Fig. 4). The proposed extended sparse KV can be represented as:
$\mathbf{K}_{\mathtt{sparse}}^{j}$ =$=\left\{K_{1}^{j},\dots,(M_{r-1}^{j}\cdot K_{r-1}^{j}),K_{r}^{j},\right.$ (4)
$\quad\left.(M_{r+1}^{j}\cdot K_{r+1}^{j}),\dots,K_{2r}^{j},\dots,K_{Z}^{j}\right\},$
$\mathbf{V}_{\mathtt{sparse}}^{j}$ =$=\left\{V_{1}^{j},\dots,(M_{r-1}^{j}\cdot V_{r-1}^{j}),V_{r}^{j},\right.$
$\quad\left.(M_{r+1}^{j}\cdot V_{r+1}^{j}),\dots,V_{2r}^{j},\dots,V_{Z}^{j}\right\}$
@@
$\mathtt{SESA}(X_{i})=\mathtt{Softmax}\left(\frac{Q_{i}K_{\mathtt{sparse}}^{T}}{\sqrt{d}}\right)V_{\mathtt{sparse}},I_{i}\in\mathbf{R}$ (5)
The proposed KV-extension in $\mathtt{SESA}(\cdot)$ selectively retains moving region tokens from all frames. Hence, slower moving regions will adaptively incorporate sparser sampling of KVs, while faster moving regions will leverage denser sampling of KVs across all sampled reference frames $\mathbf{R}$.
@@ 4.2 KV-Caching and Inter-Frame Editing
Operating extended joint self-attention across all video frames is computationally intensive. To address this, we employ sparser sampling rates of reference frames (higher $s$, $r$) in the initial phase, and perform joint editing. Subsequently, we generate intermediate frames in the second phase leveraging KV features stored from the initial phase. Previous approaches often rely on frame interpolation techniques for intermediate frames [44, 23], but these methods can result in blurring effects or inconsistent character generation. Instead of interpolation, Ada-VE focuses on directly generating intermediate frames with consistent characters to better preserve visual quality.
Our empirical observations indicate that maintaining a consistent KV feature space is crucial for generating consistent characters across intermediate frames. Rather than using only nearby KVs for intermediate frame generation [47], we propose caching the sparsely sampled KVs from all reference frames (i.e., $\mathbf{K}_{\mathtt{sparse}}$, $\mathbf{V}_{\mathtt{sparse}}$) during the first pass. These cached KVs are then reused with the queries $Q_{m}$ from all intermediate frames $I_{m}\notin\mathbf{R}$ in self-attention ($\mathtt{IFSA}(\cdot)$) during the second pass:
$\mathtt{IFSA}(X_{m})=\mathtt{Softmax}\left(\frac{Q_{m}K_{\mathtt{sparse}}^{T}}{\sqrt{d}}\right)V_{\mathtt{sparse}}$ (6)
@@ 5.1 Main results
Figure 5: Qualitative comparisons with state-of-the-art video editing methods show that SDEdit [27] struggles with motion blur and inconsistent character generation in longer videos, while ControlVideo [47] also has inconsistency issues. TokenFlow [11] offers better temporal consistency but still suffers from motion blur. In contrast, Ada-VE (ours) achieves superior visual quality and consistency throughout, building on the PnP [40] image editing baseline like TokenFlow but delivering higher-quality results.
@@
We used 15 full-length DAVIS [33] videos ranging from 40 to 200 frames, and 50 manually designed prompts to conduct quantitative evaluations based on user preference. We had 10 participants who were selected through voluntary participation. All participants were fully informed about the nature of the study. Users were asked to select the best-edited videos in three categories: (1) Visual Quality: This considers the amount of blurriness and overall visual perception of users on edited videos; (2) Temporal Consistency: This represents the temporal motion smoothness of edited videos; and (3) Content Consistency: This denotes the generation of consistent characters across all frames.
Six baseline methods were used for comparative analysis: Text2Video-zero [18], SDEdit [27], TokenFlow [11], ControlVideo [47], Motion-Transfer [45], and StreamV2V [22]. Note that only 16 frames of generated videos were considered for Text2Video-zero and MotionTransfer due to extensive memory requirements for longer videos. Moreover, we considered TokenFlow integrated SDEdit [27] to make it a stronger video baseline for a fair comparison.
Table 1: User preference study is conducted on visual quality, temporal consistency and content consistency with the same prompts. * denotes shorter video length for memory requirements.
@@
T2V-zero* [18] 1.3 1.1 0.5
MotionTransfer* [45] 12.2 15.3 14.2
SDEdit [27] 10.1 12.5 10.3
ControlVideo [47] 8.4 4.6 4.4
StreamV2V [22] 8.1 3.5 2.8
TokenFlow [11] 21.3 32.3 28.5
Ada-VE(ours) 38.6 31.2 39.3
The results are presented in Table 1. Ada-VE achieved the highest user preference in visual quality at 38.6%, enhancing visual representation with reduced blurriness, and the highest content consistency at 39.3%, generating consistent content across frames. In contrast, most other video synthesis methods struggled to generate consistent content and often exhibited temporal blurriness. Moreover, Ada-VE performed on par with the state-of-the-art TokenFlow method in terms of temporal consistency, while significantly outperforming other baseline methods. Hence, Ada-VE significantly improves visual quality and content consistency without sacrificing motion smoothness.
@@
We present a qualitative analysis of the baseline methods in Fig. 5. We highlight three strong baselines for this analysis with the same length of video: SDEdit [27], TokenFlow [11], and ControlVideo [47]. In the example of the woman running, SDEdit and TokenFlow often produce blurry fingers and hands, missing important details. Additionally, SDEdit generates inconsistent legs and distorted face structures in the moonwalk videos. TokenFlow exhibits similar face distortions. ControlVideo generates inconsistent hairstyles and hands, with more prominent inter-frame inconsistencies in the moonwalk videos. In contrast, Ada-VE generates consistent fingers in the woman running example with significantly reduced motion blurriness. Moreover, the face distortion in the moonwalk videos is notably improved in Ada-VE. These results demonstrate the superior quality of Ada-VE compared to state-of-the-art methods. Additionally, we provide extensive qualitative videos in the supplementary with all six baselines.
Figure 6: We present ablation study on video editing and style transfer prompts with proposed Ada-VE framework.
@@ 5.2 Ablation Study
We study the effects of various extensions of self-attention performance in joint editing, as shown in Table 2. We selected 40 consecutive frames from 10 DAVIS videos for joint editing based on text prompts. To quantitatively measure temporal consistency and realistic prompt alignment, we used the warping-error and CLIP-score metrics following prior work [11, 47]. Apart from the various self-attention extensions, the same baseline PnP model [40] was used for this study. Our observations based on Tab. 2 indicate that increasing the number of reference frames directly enhances performance, although it significantly increases computational burden and latency. Incorporating our proposed adaptive extension of self-attention can significantly reduce latency without compromising performance. For instance, with full extensions of self-attention, Ada-VE can achieve 3.9$\times$ speed-up on average. We use single A6000 GPUs with 48GB memory for experiments. In other words,while using the same GPU memory, proposed sparse extension facilitates use of average 3×3$\times$3 × higher reference frames than baseline full extension, with $25\%$ higher CLIP score and $45\%$ lower warp error (Tab. 3).
Table 3: With the same GPU memory, Ada-VE can operate with more reference frames contributing to performance improvement.
Extension # Ref. Frames $\uparrow$ GPU Mem $\downarrow$ Latency $\downarrow$ CLIP Score $\uparrow$ Warp Err $\downarrow$
Full Ext. 20 1x 1x 0.31 3.8
Prop. Ext. 61 1x 0.8x 0.39 2.1
@@ 7 Conclusion
In this paper, we presented Ada-VE, a novel video-to-video synthesis method designed to overcome the limitations of existing approaches. By introducing an adaptive motion-based KV selection strategy and KV-caching, Ada-VE significantly reduces computational overhead while improving visual quality and content consistency. Leveraging sparse KV feature extraction within Plug-and-Play diffusion models, our method achieves superior performance with a lower computational cost. Through extensive qualitative and quantitative evaluations, Ada-VE demonstrated substantial improvements over state-of-the-art methods, particularly in maintaining visual clarity, temporal consistency, and content fidelity across frames. Its ability to modify moving subjects consistently while preserving fine details was validated in various challenging video editing tasks. Ada-VE represents a strong advancement in video synthesis, providing a practical solution for generating high-quality, temporally consistent videos and opening new directions for future research in video editing and synthesis.
@@ Acknowledgement
This project is supported by the ONR Minerva program, iMAGiNE - the Intelligent Machine Engineering Consortium at UT Austin, and a UT Cockrell School of Engineering Doctoral Fellowship.